OpenStudy (anonymous):

Factor each trinomial x^2+2x-15

OpenStudy (anonymous):

factor (x+5) (x-3)

OpenStudy (anonymous):

x^2+2x-15 +2=(+5)+(-3) -15=(+5)(-3) therefore, x^2+2x-15 =(x+5)(x-3)